OpenStudy (anonymous):

What is the solution of the equation? Helpp! sqt2x+13 -5=x

Parth (parthkohli):

\[\sqrt{2x + 13 - 5} = x\]?

Parth (parthkohli):

Or is it\[\sqrt{2x + 13} - 5 = x\]

OpenStudy (anonymous):

The second equation! (:

Parth (parthkohli):

Okay :) Start with adding \(5\) to both sides.

OpenStudy (anonymous):

can you explain to me what sides?

Parth (parthkohli):

Add \(5\) to both sides, like this:\[\sqrt{2x +13} - 5 + 5 = x + 5\]

OpenStudy (anonymous):

ok

Parth (parthkohli):

So, what do you have now?

OpenStudy (anonymous):

\[\sqrt{2x+13}=x+5\]

Parth (parthkohli):

now square both sides :-)

OpenStudy (anonymous):

\[\sqrt{13}=5\]

OpenStudy (hba):

@kudrick16 Lets see this question :) \[\sqrt{2x+13}=x+5\] Squaring both sides

OpenStudy (hba):

\[(\sqrt{2x+13})^2=(x+5)^2\] Hint : Use (a+b)^2 formula.

OpenStudy (hba):

@kudrick16 Please solve...

OpenStudy (anonymous):

\[\sqrt{169}=25\]

OpenStudy (anonymous):

25x*

OpenStudy (hba):

@kudrick16 (a+b)^2=a^2+2ab+b^2 (x+5)^2= \[\sqrt{2x+13}=(2x+13)^{1/2}\]

OpenStudy (hba):

@kudrick16 I think you need to work on your basics first ...

OpenStudy (anonymous):

Now i don't know have a clue what too do from here on out

OpenStudy (anonymous):

well can you explain to me how to do the problem from here?